# Artworks Reimagined: Exploring Human-AI Co-Creation through Body Prompting

### v1

@@ 3.1. Background and Design Goals
The European Researchers’ Night is an annual event, which aims to interest the general public in research. It is one of the biggest science events in Finland, held at several universities nation-wide. The event gives the public the opportunity to discover the wonders of science in fun and inspiring ways. As a multidisciplinary event, Researchers’ Night offers a wide range of activities from workshops to laboratories, presentations, and lectures. The event program caters to the general public, including visiting school groups and parents with children.
Our study was born from our previous experience hosting an exhibit at the European Researchers’ Night in 2022 (Oppenlaender et al., 2023c, b). At the time, text-to-image generation was still a new phenomenon, and many visitors had not tried it before. The exhibition stand consisted of two laptops on which visitors could try out three text-to-image generation systems (Midjourney, Stable Diffusion, and DALL-E). One laptop was connected to a large public display. During this event, we observed that while interesting for the person writing the prompts, textual interaction with the image generation system was not very appealing for bystanders in a public audience, due to prompts being written in private. In the following year, we designed an interactive art installation to address this issue.
The design of our installation had three goals. First, we sought to overcome the limitations in engagement that we had observed in 2022 and provide a highly engaging experience for visitors of the local Researchers’ Night event. Second, we aimed to allow visitors to reconfigure an existing artwork with an input modality other than typing. To this end, we selected posing (“body prompting”) as input modality. Visitors would be able to reimagine an existing artwork by posing in front of a camera. Third, like a camera booth, the installation was designed to be easy to use and self-guiding, allowing its user to go through the steps necessary to take a selfie photo.
3.2. Artworks Reimagined
We created an interactive art installation, called Artworks Reimagined, that guides its users through four simple steps (depicted in Figure 1 and described in detail in Section 3.3). The physical setup of the installation is depicted in Figure 2. Divider walls were used to define four separate areas. The public posing stage allowed participants to select an artwork on the touch screen and pose in front of the camera. This setup was mirrored in the private camera booth which was cordoned off and protected from viewing with divider walls. Results from the public booth were viewed on a large screen controlled by a laptop, whereas results from the private booth were viewed in private on two laptops equipped with privacy screens.
We selected artwork images for the installation from two sources. The first source included paintings from the “Golden Age of Finnish Art” (circa 1880-1910). We identified artworks by visiting Wikipedia pages of prominent artists from that time period. [^2] https://en.wikipedia.org/wiki/Golden_Age_of_Finnish_Art The second source was a list of popular paintings on WikiArt. [^3] https://www.wikiart.org/en/popular-paintings/alltime When selecting artworks, we avoided nudity and other depictions not appropriate for public viewing in the presence of children. This convenience sample of 60 artworks (29 from the Finnish Golden Age and 31 from WikiArt) represents a mix of well-known international masterpieces and artworks with a deep local context.
@@ 3.2. Artworks Reimagined
The study setup consisted of four separate areas: 1) the public viewing and interview area, 2) the public posing stage with public display, 3) the private posing area with display, and 4) the private viewing area with two laptops equipped with privacy screens.
3.3. Study Design and Procedure
The aim of our study was to observe and understand how body prompting is used and experienced in the context of a public event. Therefore, the research was conducted “in-the-wild”, focusing on observing participants in an environment outside of the controlled laboratory, without interference from researchers. The research was exploratory and participants were not given specific posing instructions.
The study procedure is summarized in Figure 1. Participants were invited to walk up to the public display. The first application screen consisted of a consent form informing the participants about the extent and use of the collected data. Participants were specifically informed that only an anonymous, stick figure-like pose would be collected, not the original photo. After providing consent by clicking a button, participants selected an artwork from the list by touching the screen. Next, participants were informed that a timer would count down from 10 seconds. After confirming this message, participants could see themselves on the screen and assume a pose while the timer counted down to zero. The posing was supported by showing the artwork for a brief moment before displaying the camera feed, and a small version of the artwork was additionally shown in the corner of the screen while posing. Participants were not given any instructions on what specific pose to assume and were completely free to deviate from the source artwork. Finally, participants were shown a short unique code from a curated list of Finnish and English words, generated with a language model, and informed to exit the booth in the final application screen. The short code was used to associate the subsequent interview data with the generated image.
Following the interaction with the installation, one of five research assistants invited the participants to a semi-structured interview recorded using a mobile phone. The assistants showed the participant(s) the artwork and detected body prompt during the interview and recorded participant(s)’ reactions and comments in response to seeing the generated artwork.
Figure 3. Application architecture with inputs (left), detected textual prompt and body prompt (middle), and output (right).

### v2

@@ 3.1. Formative Study
Our work was informed by our prior experience of hosting an exhibit at the European Researchers’ Night in 2022. The European Researchers’ Night is an initiative funded by the European Union under the Marie Skłodowska-Curie Actions (MSCA). It is the largest science outreach event in Europe, aiming to bring researchers closer to the public and showcasing the impact of research on daily life. The event takes place annually at universities in numerous European countries. The European Researchers’ Night aims to interest citizens in research. To this end, the event gives the public the opportunity to discover the wonders of science in fun and inspiring ways. As a multidisciplinary event, Researchers’ Night offers a wide range of activities from workshops to laboratories, presentations, lectures, and exhibits. The event program caters to the general public, including visiting school groups and parents with young children.
Text-to-image generation was still a novel phenomenon in 2022. We decided to demonstrate this novel technology to visitors of the European Researchers’ Night. Our exhibition stand at the event allowed visitors to try out three existing text-to-image generation systems (Midjourney, Stable Diffusion, and DALL-E). The exhibition stand consisted of two cocktail tables with one laptop each. One of the laptops was connected to a large public display, the other allowed image generation more privately without displaying results publically.
During this well-visited event, we had numerous conversations with visitors, noting that many visitors had not tried text-to-image generation before. One central insight from our observations and conversations with visitors was that, while interesting for the person writing the prompts, text-based interaction with the image generation system was not very appealing and engaging for an audience in a public setting. While prompts were being written in private, and users were deeply engaged in writing these prompts, other visitors could simply not see what was going on. The interaction with the generative AI was hidden from the public eye. For passers-by, the mechanics of the public display installation were not obvious. That is, it was not immediately obvious that prompts written in private on the laptop manifested as images on the large screen.
For the following year, we built on this important observation and designed an interactive art installation to address this issue. In particular, we sought to address the “air gap” between input and screen, to provide a more interactive and expressive experience to the users, and a more engaging experience for all other visitors. In the following section, we summarize our formative study’s findings by outlining four key challenges that limit the application of text-to-image generation in public event settings.
3.2. Challenges of Text-to-Image Generation for Public Event Settings
From the formative study, we distill four key challenges of text-based image generation that limit its use in public event settings: output control, literacy, collaboration & social creativity, and expressivity.
  • C1.
    Output control: The level of control over text-to-image generation is limited by language (McCormack et al., 2023). One cannot describe an image with discrete prompts in every detail, leaving much of the initial image generation to randomness (Hauhio et al., 2023). “Prompt engineering” (Oppenlaender, 2024, 2022, 2023; Oppenlaender et al., 2024) for text-to-image generation is tedious, and controlling the image generation process requires a considerable amount of practice and skill (Oppenlaender et al., 2024).
  • C2.
    Literacy: Frequently regarded as an instrumentalist use of technology, text-to-image generation is a process in which AI is used as a tool to visually translate concepts or ideas (Coeckelbergh, 2023). For this translation to be effective, English literacy is required. Further, some image generators require the use of specific keywords to produce high-quality outputs (Oppenlaender, 2023), which demands expertise in “prompt engineering” and familiarity with art concepts, such as styles and media (Oppenlaender, 2023). These language-based requirements create barriers of accessibility and exclude certain populations—such as the illiterate, young children, and all those who are not familiar with the English language—from participating.
  • C3.
    Collaboration & social creativity: Contemporary literature views creativity as a social activity (Frich et al., 2019; Amabile, 1983), with the concept of the lone creator being dispelled as a myth (Lemley, 2012). Yet, image production with generative AI is almost exclusively completed at home by a solitary creator. The social and performative aspects of art creation are lost due to the privacy of the image generation setting. The text input modality is not very engaging for an audience in a public event setting. Collaborative websites, such as Artbreeder [^1] https://www.artbreeder.com, have demonstrated that collaborative image generation can be an engaging experience, with wide audience appeal.
  • C4.
    Expressivity: Confined to the narrow triangular interaction space between user, keyboard, and computer display, the practice of text-to-image generation does not lend itself well to public settings, such as Galleries, Libraries, Archives, and Museums (GLAM). For such public settings, input modalities should be dynamic and intuitive, offering glimpses of the generative process to the audience and building anticipation for the final image. For a public event setting, engaging larger groups of people beyond the individual, is required. Typing on a keyboard in front of a computer screen is not expressive, in this regard, and does not provide an engaging experience to an audience.
These challenges, while not being insurmountable in the use of text-to-image generation by an individual practitioner from home, present challenges to the use of text-to-image generation technology in installations at public event settings. In the following sections, we explain the design goals for our interactive art installation (Section 3.3), and how we aimed to address these challenges in the design of our installation (Section 4).
3.3. Design Goals
In our design of the interactive art installation, we sought to overcome the limitations in engagement that we had observed at the prior European Researchers’ Night event. Our aim was to provide a highly engaging experience for visitors of this event. Based on our formative study and the four challenges delineated above, we formulate the following five design goals for our art installation in a public event setting:
  • D1
    Enhance user control over image generation: By allowing body movements to guide image creation, the installation gives users a direct and intuitive way to shape outcomes. This design goal addresses the challenge of output control (C1) by reducing reliance on imprecise text prompts and enabling more immediate and expressive creative decisions (C4).
  • D2
    Reduce language dependency: Leveraging body prompting removes the need for specialized textual skills or prompt engineering (C2). This makes the experience more inclusive by lowering language and literacy barriers, ensuring that people of different ages, language proficiencies, and technical backgrounds can participate.
  • D3
    Facilitate social creativity and collaboration: The system is designed to encourage group participation and shared creative exploration (C3). By shifting from solitary, text-based input to embodied interaction, it supports collaborative artistic expression and reintroduces the social dimension of art-making, which is often lost in generative AI setups.
  • D4
    Support dynamic and expressive interactions: Public installations benefit from interactions that are visibly engaging and performative. This goal ensures that the system’s real-time feedback and visual display supports the posing process and creates an engaging experience for both active participants and observers, addressing the expressivity challenge (C4) inherent in static text-based interfaces.
  • D5
    Accommodate varied interaction contexts: Recognizing that some users may feel inhibited in public performance, the installation provides both public staging and private booth settings. This design goal aims to offer flexibility so that users can choose the interaction context that best suits their comfort level, thereby fostering participation across a diverse audience.
The design goals outlined above provided a framework for addressing the challenges of text-based image generation in public installations. We now turn to the system design and implementation, detailing how our architecture and technical choices align with the design goals. In the following section, we describe the system’s overall concept, its modular pipeline and key components, and interaction mechanisms that enable embodied co-creation in dynamic public settings.
4. Artworks Reimagined
We designed and developed an interactive art installation, called Artworks Reimagined, that allows its users to engage with image generation technology with a low barrier of entry.
4.1. Overall Concept Addressing the Design Goals
Artworks Reimagined is an interactive installation that leverages “body prompting” to enable users to reimagine classical artworks through their own physical expressions. Rather than relying on traditional text-based inputs, visitors use their body poses to guide a generative AI, resulting in transformations of existing digital artwork images. This approach reconnects the creative process with embodied art-making, offering a more natural and expressive means for users to interact with technology. The installation was designed for public events, featuring both an open staging area and a private booth to accommodate different comfort levels. By integrating pose detection with a generative pipeline, Artworks Reimagined invites visitors to engage in a collaborative and dynamic reinterpretation of art, bridging the gap between traditional creative practices and contemporary AI technologies.
4.2. Materials
As source materials for the interactive art installation, we aimed to select materials that would be interesting and engaging to visitors. To this end, we selected digital artworks from two sources.
The first source included paintings from the Golden Age of Finnish Art (circa 1880-1910). The Golden Age of Finnish Art is central to Finnish identity, as it represents a transformative era when artists celebrated Finland’s unique natural beauty, cultural heritage, and emerging national spirit, thereby reinforcing a strong sense of pride and self-determination. We identified artworks from this era by visiting Wikipedia pages of prominent artists from that time period. [^2] See https://en.wikipedia.org/wiki/Golden_Age_of_Finnish_Art
The second source was a list of popular paintings on WikiArt. [^3] See https://www.wikiart.org/en/popular-paintings/alltime We purposefully selected artworks from this list, taking care to avoid including nudity and other depictions not appropriate for public viewing in the presence of children.
This convenience sample of 60 artworks (29 from the Finnish Golden Age and 31 from WikiArt) represents a mix of well-known international masterpieces and artworks with a deep local context.
4.3. Technical Implementation
We developed a software architecture (see Figure 1) enabling us to generate images based on four inputs: the user’s pose, the existing artwork, and two prompts. This architecture allows us to generate a digital image that resembles the style of the original artwork, but is adapted to a different pose. In Section 4.3.2, we provide details on the implementation of our system’s backend, where images are created via API. This is followed by a description of the frontend (user interface) in Section 4.3.3.
Figure 1. Application architecture with inputs (left), textual prompts and body prompt (middle), and output (right).
4.3.1. Output control
In the context of image generation with generative AI, ControlNet (Zhang et al., 2023) was a remarkable breakthrough, allowing fine-grained control over image generation via human poses and other input types (such as depth masks, edge maps, and line drawings). A pose in ControlNet is a skeleton-like structure with key points that correspond to major joints and body parts. Our installation uses ControlNet to translate captured body poses into structured constraints that guide the image synthesis process. Specifically, a real-time camera feed allows users to assume a specific pose. A photo of the user’s pose is taken from this stream, and a pose detection algorithm (OpenPose (Cao et al., 2021)) extracts key points to form a skeletal representation of the user’s body prompt. This representation is then provided as input to ControlNet, which generates an image layout aligned with the user’s pose. the generative processing also integrates style transfer techniques and automated prompt refinement to balance user control with algorithmic creativity. This layered approach ensures that the final output reflects both the expressive intent of the user’s movements and the inherent unpredictability of generative AI, thereby addressing the challenge of output control in public art installations (C1).
4.3.2. Backend for image generation
The central component in our architecture is T2I-Adapter (Mou et al., 2023). T2I-Adapter is a controllable image generation framework that extends text-to-image (T2I) diffusion models by incorporating additional structural guidance, such as edges, depth, or keypoints, to better align generated images with user inputs. “Adapters” are used by T2I-adapter to receive the inputs. In our case, T2I-adapter receives four different inputs, as follows (cf. Figure 1):
  1. (1)
    Pose: A webcam image took a photo of the user. The user’s pose was then detected from this body prompt photo with openpose-based ControlNet (Zhang et al., 2023; Cao et al., 2021). In our system, the pose did not include facial expression and hand movements for improved anonymity. The pose was then used to guide the image generation via T2I-Adapter’s pose adapter.
  2. (2)
    Artwork: Style information was extracted from the source artwork and used as an input for the T2I-Adapter, using the Style Adapter in T2I-Adapter. The Style Adapter is a module that guides text-to-image diffusion models to generate images with a specific artistic or visual style by conditioning the generation process on reference style features.
  3. (3)
    Prompt: A textual prompt was generated from the selected source artwork using CLIP-Interrogator (version 0.6.0). [^4] https://github.com/pharmapsychotic/clip-interrogator CLIP-Interrogator is a tool that uses OpenAI’s CLIP (Radford et al., 2021) model to analyze an image and generate a descriptive text prompt that could have produced it. This prompt was given to T2I-Adapter to guide the image generation, in addition to the other inputs.
  4. (4)
    Negative prompt: A negative textual prompt was manually designed, based on best practices in text-to-image generation (Oppenlaender, 2023). Negative prompts are commonly used to mitigate generative glitches, prevent certain subjects from appearing, and to avoid low quality outputs. The same negative prompt was used for all image generations to improve the quality of outputs and prevent unsafe generations (Qu et al., 2023).
Within T2I-Adapter, images were generated with Stable Diffusion version 1.5 (Rombach et al., 2022) with 50 diffusion steps and CFG 8.0. The CFG (Classifier-Free Guidance) parameter in text-to-image generation controls the balance between faithfulness to the prompt and diversity in the generated images. The generated images were upscaled twice with Real-ESRGAN (Wang et al., 2021b) to increase the images’ resolution for viewing on the large public display. GFPGAN (Wang et al., 2021a) was used to enhance faces. The image generation architecture was packaged as a cog container (Replicate, 2024) and uploaded to Replicate.com where it is publicly available via a user interface and an API. [^5] https://replicate.com/joetm/camerabooth-openpose-style
4.3.3. Frontends for guiding the process and viewing the results
Two separate web applications were developed, one for guiding the body prompting process and one for viewing the results. We decided on two applications because we wanted to avoid users from blocking the body prompting stage when viewing results. Instead, users viewed the results on a separate display (cf. Figure 3).
Both applications were created with React, a JavaScript web development framework, and hosted as static websites in AWS S3 buckets, using an ssl-encrypted AWS Cloudfront to provide a secure connection. The secure connection was required by the MediaStream Web API used for acquiring images from the webcam. Much of our installation’s functionality will become clearer in our description of the user study in Section 5. In the remainder of this section, we describe the technical implementation of the two applications.
Web Application 1: Body prompting
The first application guided users through the process with a linear application flow (c.f. Figure 2). After providing consent, a timer was started on the click of a button. Once the timer hit zero, the application captured the photo from the webcam MediaStream using the MediaDevices API. The processing of the photo was then completed in the cloud, by sending an API request with the image as payload (in binary format) to Replicate’s servers. Replicate [^6] https://www.replicate.com is a platform that enables users to run machine learning models in the cloud via a simple API, providing access to various AI models without requiring local setup.
An AWS Gateway was set up to proxy API requests from our web application to Replicate’s API. This was done to avoid issues with CORS (cross-origin resource sharing) in our web application. Once completed, results from Replicate’s API were received with a webhook implemented as an AWS Lambda function and stored in a JSON file hosted in an AWS S3 bucket. We also stored log files of all interactions with the system in a separate AWS S3 folder.
Note that to maintain data privacy, the original images were not stored. Instead, only the generated image and the pose was stored for further analysis. This was clearly communicated to the user in the consent form (which was the first screen in the application – see Section 5).
After a picture was taken and sent to the server for processing, the application was automatically reset once 60 seconds had passed. The main purpose of this timed reset was to reduce queuing for participants at the public viewing area—the bottleneck of our installation and study design—and to give more time to the subsequent interviews. The source artworks were randomly shuffled after each image generation to provide each participant with an unbiased list of artworks to select from.
Web Application 2: Image viewing
The second application was a light-weight frontend created for viewing the resulting images. The application used long-polling to fetch the list of generated images stored as JSON file in an AWS S3 bucket. The research assistants could scroll back and forth through the images with keyboard arrow keys. A red square on the bottom right of the screen alerted the assistant that newly generated images were available for viewing. The number of new images was displayed in this red square. The participant’s pose was shown in the bottom left corner of the screen. Generated images were viewed either on a large display in public or in private on two laptops with privacy screens, depending on which posing area was chosen by the participant.
5. Evaluation
The aim of our study was to evaluate body prompting as a novel input modality for image generation in public settings. We aimed to observe and understand how body prompting is used and experienced in the context of a well-visited public event. Therefore, the research was conducted “in-the-wild”, focusing on observing participants in an environment outside of the controlled laboratory, without interference from researchers. The research was exploratory and participants were not given specific posing instructions.
5.1. Study Procedure
The study procedure is summarized in Figure 2 and the spatial layout is depicted in Figure 3. Participants were invited to walk up to the public display. The first application screen consisted of a consent form (①) informing the participants about the extent and use of the collected data. Participants were specifically informed that only an anonymous, stick figure-like pose would be collected, not the original photo. After providing consent by clicking a button, participants could scroll through the artworks, which were presented in a gallery with masonry layout (②). Participants selected an artwork from the gallery by touching the screen. Next, participants were informed that a timer would count down from 10 seconds. After confirming this message, participants could see themselves on the screen and assume a pose while the timer counted down to zero (③). The posing was supported by showing the artwork for a brief moment before displaying the camera feed, and a small version of the artwork was additionally shown in the corner of the screen while posing. Participants were not given any instructions on what specific pose to assume and were completely free to deviate from the source artwork. Finally, participants were shown a short unique code from a curated list of Finnish and English words, generated with a language model, and informed to exit the booth in the final application screen (④). The codes were carefully reviewed to avoid introducing bias into the study. The short code was used to link the subsequent interview with the generated image.
Figure 2. The study setup of the interactive art installation, Artworks Reimagined, guides participants through four application screens (top) in which participants give consent ①, select an artwork ②, body prompt via posing in front of a camera ③, and exit the camera booth ④. This is followed by an interview in which the results are revealed ⑤.
[…]
Following the interaction with the installation, one of five research assistants invited the participants to a semi-structured interview recorded using a mobile phone (⑤). The assistants showed the participant(s) the artwork and detected body prompt during the interview and recorded participant(s)’ reactions and comments in response to seeing the generated artwork.
@@ 5.1. Study Procedure
The study was conducted in accordance with the policies of the University of Jyväskylä’s Institutional Review Board (IRB) and the Finnish National Board on Research Integrity (TENK). Children under the age of 18 were allowed to use the installation under the supervision of adults, but were excluded from the interviews. In this article, we only focus on the data collected from 79 adults who agreed to an interview following the usage of the installation. The semi-structured interview questionnaire is described in the following section.
5.2. Data Collection and Analysis
We developed an interview guide consisting of 12 questions. Besides basic demographics, the research assistants asked participants about their motivation for selecting the artwork, the body prompt, and the choice of private or public posing area. The interviewers also inquired how participants experienced body prompting and what they wanted to express with their body prompt. Some questions were followed by open-ended questions (“can you describe why?”). A Likert-scale question captured how enjoyable the experience was for the participant (from 1 – Not At All Pleasant to 5 – Extremely Pleasant). The interview also included the ten-item questionnaire on key personality traits, commonly referred to as “Big Five” (Digman, 1990).
The generated image was revealed on the screen at the end of the interview and participants were asked to comment on how they perceived the image and whether the image expressed their original intent, thus giving insights into co-creative aspects of body prompting, and any other thoughts about their body prompt and the generated image. Last, participants were asked to rate and comment on the overall experience (on a Likert scale from 1 – Not At All Enjoyable to 5 – Extremely Enjoyable).
The interview data was transcribed using a text-to-speech service, translated from Finnish into English using Google Translate, and reviewed and corrected by three research assistants who listened to the recordings and manually corrected translation errors. The qualitative data was analyzed following a grounded theory approach (Glaser and Strauss, 1967). Three authors (one Postdoc in Cognitive Science, one PostDoc in Computer Science, and one PhD student with professional experience in user experience design) discussed a preliminary coding scheme for each question as a starting point for inductive coding. The three authors then individually coded the responses of a sample of 30 participants. In a discussion following this individual coding session, the authors noticed that coding the data was straightforward (McDonald et al., 2019). That means, in all but extremely few cases, it was easy to identify themes in the data, and in two other questionnaire items, the responses were too diverse to identify common themes. Therefore, it was decided that one author would inductively code the interview data, paying attention to internal consistency of the data coding. Results were then discussed in two meetings.
Further, the generated images and body prompts (i.e., poses) were rated by three authors along the following criteria: a) dynamism of the source artwork (low/medium/high), b) dynamism of the participant’s pose (low/medium/high), and c) a comparison of the source artwork and generated image in terms of whether a change in narrative took place. For a) and b), the authors collaboratively agreed on a codebook, as listed in Table 2 in Appendix A. With the metric of ‘dynamism’, we intend to capture the expressiveness of the artwork and body prompt. A static pose was rated to be of low expressiveness, whereas a highly dynamic pose was rated as being very expressive. For c), the authors agreed to code a change in narrative if a subject engaged in a different activity or had a different expression (e.g., static versus humorous), or if the model hallucinated artifacts that caused or strongly contributed to a change in narrative between the source image and the generated image. After a first rating round, the ratings were discussed and each author revised their ratings for internal consistency.
Employing Fleiss’ Kappa for inter-rater reliability assessment, our study found high agreement levels among the three raters across all three variables: source artwork dynamism ($\kappa=0.96$), pose dynamism ($\kappa=0.86$), and narrative change ($\kappa=0.88$). These results underscore the raters’ cohesive interpretations and a consistent and robust evaluation process for the studied variables. A majority vote between the three authors was used to produce the final set of ratings, in cases where the authors’ ratings diverged.
In addition, one author analyzed and coded the difference between the source artwork and the body prompt, characterizing the latter as either being an imitation (re-creation) of the source artwork or an attempt to ‘reimagine’ (reinterpret) the artwork with a body prompt that significantly differs from the source artwork. The latter reflects the participant’s intention to shift the narrative between source artwork and generated output by using a body prompt that introduces significant differences. It was further coded whether this narrative change in the resulting image originated from the generative model (e.g., by hallucinating artifacts or facial expressions) or the user (via the body prompt), with options ‘model’, ‘user’, ‘both’, and ‘n/a’ (meaning no major change in narrative took place).
5.3. Participants
Interviews were conducted with 79 visitors (henceforth called participants) after they used the installation. Participants were on average 37 years old ($Median=35$ years, $SD=13.5$ years, $Max=68$ years) and included 46 women, 32 men, and one non-binary participant.
Due to the study being conducted in the field without control over the number of participants per condition (i.e., group versus solo participation, body prompting strategy, artwork choice, and public versus private posing area), we focus on describing the images in Section 5.4.1 and qualitatively analyzing the interview data in sections 5.4.25.4.5. For a rich description, we contextualize participant quotes with the codes listed in Table 1.
Table 1. Contextual information on participants.